The rate of effusion of NH3 gas through a porous barrier is observed to be 1.15*10^-4 mol/h.

Under the same conditions, the rate of effusion of N2 gas would be _________
mol/h.
1
Expert's answer
2020-06-08T15:30:43-0400

As per Graham's law,

Rate of diffusion /effusion is inversely proportional to the molar mass of the gas.

And "M_{NH_3}=17\ g; M_{N_2}=28\ g"

"\frac{r_{N_2}}{r_{NH_3}}=\sqrt{\frac{M_{NH_3}}{M_{N_2}}}" "\implies" "\frac{r_{N_2}}{1.15\times10^{-4}}=\sqrt{\frac{17}{28}}" "\implies r_{N_2}=8.97\times 10^{-5} mol/h"